OpenStudy (anonymous):

dealing with porabolas and quadratic equations... i need help with this problem , y= -x^2 -3x =18 not sure if im doing it right, but when i tried solving for y, & to solve for y you would use -b/2a , i got 3/2.. am i right?

OpenStudy (baldymcgee6):

the equation already has y isolated... -b/2a is to find the vertex

OpenStudy (anonymous):

oops the equation is y=-x^2-3x+18

OpenStudy (anonymous):

tht is what im trying to find, the vertex to plot the porabola

OpenStudy (baldymcgee6):

okay, then you are on the right track with that formula, let me check...

OpenStudy (baldymcgee6):

You should find the vertex to be at (-3/2, 20.25)

OpenStudy (baldymcgee6):

Once you use the vertex formula to find the x value of the vertex, plug it back in to fund the y value

OpenStudy (baldymcgee6):

-b/2a = -2/3 y=-(-2/3)^2-3(-2/3)+18 y = 20.25

OpenStudy (anonymous):

so it should be a fraction. and yep thats what i got thanks.!

OpenStudy (baldymcgee6):

you're welcome